assume the inflation rate 6% per year and the real interest rate is 5%.

A)the number of future dollars after 5 years that will be equivalent to 30000 calculating first real dollar equivalence (constant value dollars)

B) The number of future dollars after 5 years that will be equivalent 30000 market interest rate ?
Answer #1

A) FV = PV * (1+inflation rate)^number of years

FV = 30,000 * (1+0.06)^5

FV = 30,000 * 1.06^5

FV= 40,146.77

B)

FV = PV * (1+ Real interest rate)^number of years

FV = 30,000 * (1+0.05)^5

FV = 30,000 * 1.05^5

FV= 38,288.45